# Improving the Efficiency of Self-Supervised Adversarial Training through Latent Clustering-Based Selection

### v1

@@ 6.2. Hyperparameter Sensitivity
Ratio of Selected Unlabeled Data $\bm{\alpha}$. Figures LABEL:fig:cofarper and LABEL:fig:svhnper illustrate how the performance varies with selection ratio $\alpha$ for CIFAR-10 and SVHN, respectively. We analyze how the quantity of extra data influences the effectiveness of our data-selection method, which is designed to prioritize data points located near decision boundaries. When the external data supply is limited, our algorithm excels by selecting a higher proportion of boundary-adjacent data, achieving a balanced selection that enhances robust accuracy. However, as the volume of external or generated data increases, the algorithm’s focus on boundary points can inadvertently lead to overfitting, diminishing the improvement gained on the model’s generalization ability. Interestingly, as the data volume continues to grow and more points are added, the performance begins to improve once again, suggesting that the initial overfitting is mitigated by the sheer abundance of data, leading to a recovery in accuracy.
Table 2. Performance of intermediate models trained using different schemes, including self-supervised learning (SSL), pre-trained CLIP, and fully-supervised methods, on CIFAR-10 under varying ratios of labeled data. We also present the corresponding performance of the final model produced by SSAT with LCS-KM, where the ratio of selected unlabeled data $\alpha$ is set at $10\%$.
Labeled Ratio Intermediate Model Final Model (LCS-KM, $\alpha=10\%$)
[…]
$1\%$ SSL ($5\%$ unlabeled) $78.8$ $82.9$ $39.8$
pre-trained CLIP $86.8$ $83.9$ $40.5$
$10\%$ SSL ($5\%$ unlabeled) $82.9$ $85.0$ $42.4$
pre-trained CLIP $86.8$ $85.3$ $43.2$
$20\%$ SSL ($5\%$ unlabeled) $87.0$ $87.2$ $48.3$
pre-trained CLIP $86.8$ $85.6$ $47.7$
100% fully-supervised $90.5$ $86.2$ $58.2$
pre-trained CLIP $86.8$ $86.5$ $57.4$
6.3. Further Exploration of Generalizability
All experiments in this section are conducted on CIFAR-10. We compare four different settings: random selection with $\alpha$ ratio of unlabeled data, selecting $\alpha$ ratio using LCS-KM, using only labeled data (i.e., $\alpha=0\%$), and using the entire dataset (i.e., $\alpha=100\%$).
Perturbation Size. We consider external data from TinyImages as the unlabeled dataset. We investigate the effect of varying the $\ell_{\infty}$ perturbation size $\epsilon$, setting it to $4/255$ and $16/255$ while keeping TRADES as the adversarial training algorithm similar to the experiments in Table 1. The results are depicted in Figures LABEL:fig:4_255 and LABEL:fig:16_255, suggesting that our LCS-KM method is highly effective compared with random selection and no selection, almost approaching the robust accuracies achieved by full utilization of unlabeled data. The trends are shown to be consistent across varying perturbation sizes.
Adversarial Training. We explore the impact of replacing TRADES with vanilla AT as the training algorithm. Similarly, we use external data from TinyImages as the unlabeled dataset. The results for PGD-based AT are shown in Figure LABEL:fig:-pgd. Additionally, we reduce the number of PGD steps from $10$ to $5$, and the outcomes are summarized in Figure LABEL:fig:pgd-5. Across all configurations, our LCS-KM selection consistently outperforms random selection and achieves accuracy comparable to using the entire labeled dataset, suggesting the generalizability and robustness of our methods.
$\bm{\ell}_{2}$ Perturbations. We further evaluate the performance of LCS-KM under $\ell_{2}$ perturbations with size $\epsilon=128/255$. In this experiment, we augment CIFAR-10 with either external unlabeled data sourced from Tiny-ImageNet or synthetic data generated via DDPM. We consider three configurations of selecting $10\%$, $20\%$, and $100\%$ extra unlabeled data. Our findings are shown in LABEL:fig:externall2 and LABEL:fig:generatedl2, revealing that selecting $20\%$ unlabeled data achieves nearly equivalent robust accuracy compared to using the full dataset with $\alpha=100\%$. These results again confirm the efficiency and effectiveness of LCS-KM when adopted in self-supervised adversarial training algorithms.
6.4. Low Labeled Data Regime
We further study how the choice of the intermediate model and the amount of labeled data impact the final performance. For all experiments, we use a subset of the CIFAR-10 dataset as labeled data, supplemented with 10% of data selected from a pool of $500$K unlabeled images sourced from Tiny-ImageNet. This additional data is selected using the intermediate model. The amount of unlabeled data remains constant across experiments, while the quantity of labeled data is varied. The intermediate model is trained using self-supervised learning (SSL) with a $5\%$ randomly selected subset of the unlabeled dataset, as long as the labeled data ratio is low. Additionally, we include a pre-trained CLIP model (Radford et al., 2021) as an alternative intermediate model for comparison. The results of these experiments are presented in Table 2. The intermediate model’s quality significantly impacts the final model’s robust accuracy. In low-label regimes (e.g., $1\%$ or $10\%$), models trained using our intermediate model yield comparatively lower robust accuracy than those leveraging the pre-trained CLIP model. However, as the amount of labeled data for the intermediate model increases (e.g., 20% or 100%), the final model results are similar to or better than those of its counterparts based on the CLIP intermediate model. This trend demonstrates the effect of the intermediate model on the final robust accuracy. These observations underscore the importance of having a strong intermediate model for robust performance. Additionally, we find that pre-trained models like CLIP can be a good alternative to training an intermediate model, especially in low-label regimes. Detailed experiment results and discussions on the pre-trained CLIP intermediate model are provided in Appendix D.3.
6.5. Further Discussion
Our study introduces strategic data selection methods, including Latent Clustering-Based Selection with k-means (LCS-KM), Latent Clustering-Based Selection with Gaussian Mixture Models (LCS-GMM), and prediction confidence-based selection, to improve the efficiency of self-supervised adversarial training. These approaches tackle key challenges in adversarially robust learning, particularly the significant memory and computational overhead of existing methods. By prioritizing boundary-adjacent data, these processes aim to enhance model robustness. However, these methods assume that the model’s latent representations are reliable for clustering and confidence evaluation, an assumption that may falter in the case of undertrained or poorly calibrated models. While these approaches effectively reduce the volume of unlabeled data required for training, their performance can be sensitive to hyperparameter configurations, such as the balance between boundary and non-boundary points and the size of the unlabeled dataset. By focusing on challenging examples near the decision boundary, these selection methods naturally address some aspects of robust overfitting and support robust generalization. Nevertheless, deeper investigation is required to fully comprehend how these strategies influence decision boundary geometry and the model’s capacity for robust generalization. Combining these selection techniques with fast and efficient adversarial example generation methods, such as fast and free adversarial training, holds promise for further accelerating the training process while preserving robust performance.
7. Conclusion
In this work, we illustrated the importance of latent clustering-based data selection in improving the efficiency of SSAT while maintaining robust accuracy. Emphasizing boundary data points during selection significantly enhances the model’s overall robustness. For SVHN, we also showed improvement in robust accuracy by adding ground-truth labels to a small set of data. This suggests manually annotating a few unlabeled data points close to the boundary could be a potential future direction to further enhance model robustness with considerations of memory and computational efficiency. Our work opens new avenues for further research, including exploring advanced data selection schemes, making the selection process more interpretable, and investigating the trade-offs between model performance, data selection, and computational costs.
Appendix A Overall Training Pipeline
Figure 6 illustrates the overall pipeline, based on the robust self-training framework proposed by Carmon et al. (Carmon et al., 2019). We first train our intermediate model using the available labeled data and use this intermediate model to select a subsection of our unlabelled or generated data. The final robust training is performed on all the labeled data and a strategically selected subset of the unlabelled or generated data.
Figure 6. Overview of the training pipeline. An intermediate model is first trained with labeled data and then used to select data from the unlabeled set. Finally, we adopt SSAT to learn a robust model on both the labeled and the reduced unlabeled datasets.
@@ Appendix B Detailed Experimental Setup
SVHN. SVHN is naturally divided into a core training set comprising approximately $73$K images and an extra training set of around $531$K images. Initially, the model is trained on the $73$K labeled images. To evaluate the generalizability of our selection schemes, we apply it to a synthetic dataset generated using a DDPM model (Gowal et al., 2021).
@@
Configuration. For model training, we adopt a WideResNet 16-8 (WRN-16-8) architecture (Zagoruyko and Komodakis, 2016) in all our SVHN experiments. We generate adversarial examples using PGD attacks exactly as implemented by Zhang et al. (Zhang et al., 2019), with step size $0.007$, $10$-step PGD attack iterations, and $\ell_{\infty}$ perturbation magnitude $\epsilon=0.015$. Hyperparameters are set the same as in Carmon et al. (Carmon et al., 2019) except for the number of epochs. For SVHN, we use a training batch size of $128$. All of our experiments that utilize all of the extra 531K data are run for $200$ epochs and the experiments with $1$M generated data are run for $400$ epochs. To prevent overfitting and reduce computational complexity, we run our experiments with selected data for $75$ epochs with early stopping. Following prior literature (Carmon et al., 2019), we adopt an SGD optimizer with a weight decay factor of $5\cdot 10^{4}$ and an initial learning rate of $0.1$ with a cosine learning rate annealing to train the model.
@@
Evaluation. For the attack evaluation to calculate the robust accuracy using PGD attack, we keep the parameters similar to that of Carmon et al. (Carmon et al., 2019) for better comparison. We use step size $0.005$, number of attack steps $K=100$. We evaluate models at $\epsilon=0.015$, which is the same as the value we used during training. We also use Auto Attack with the same $\epsilon$ value as in PGD attack.
@@
CIFAR-10. The CIFAR-10 dataset has 50K labeled images. We use the $80$ Million Tiny Images (80M-TI) dataset, of which CIFAR-10 is a manually labeled subset, to obtain extra data. However, most images in $80$M-TI do not correspond to CIFAR-10 image categories. Carmon et al. (Carmon et al., 2019) used an 11-way classifier to distinguish CIFAR-10 classes and an 11th “non-CIFAR-10” class using a WideResNet 28-10 model. Each class selected an additional $50$K images from 80M-TI using the model’s predicted scores to create a $500$K pseudo-labeled dataset, which we used in our experiments. We train the intermediate model with $50$K labeled data and use this model to select data from the $500$K pseudo-labeled data or a synthetic dataset. We perform the same experiments as in SVHN, except that we do not have the ground-truth labels of the additional unlabeled dataset. We also conducted further experiments utilizing $1$M synthetical data generated by the DDPM model (Gowal et al., 2021).
@@
Configuration. For our main CIFAR-10 experiments, we use a WideResNet 28-10 (WRN-28-10) (Zagoruyko and Komodakis, 2016) architecture to train models. Similar to SVHN experiments, we set step size $0.007$, PGD attack iterations as $10$ and $\ell_{\infty}$ perturbation magnitude $\epsilon=0.031$. For our adversarial training with $\ell_{2}$ perturbation, we use $\epsilon=128/255$. In section 6.3 when $\epsilon=4/255$, we use $0.0035$ as the PGD step size. For $\epsilon=16/255$, we use $0.015$ as the PGD step size. Hyperparameters used are the same as in Carmon et al. (Carmon et al., 2019) except for the number of epochs. Here, we use a training batch size of $256$. The experiments using all the $500$K pseudo-labeled data are run for $200$ epochs, and using 1M generated data are run for $400$ epochs. For our experiments with limited data, we get the best results at $100$ epochs where we early stop the training process. Following prior literature (Carmon et al., 2019), we adopt an SGD optimizer with a weight decay factor of $5\cdot 10^{4}$ and an initial learning rate of $0.1$ with a cosine learning rate annealing to train the model.
@@
Evaluation. The PGD attack evaluation is conducted similarly to that of Carmon et al. (Carmon et al., 2019) for fair comparisons. We use step size $\alpha=0.01$, number of attack steps $K=40$. We consider $\ell_{\infty}$ perturbations with $\epsilon=0.031$, the same as the perturbation magnitude we used during training. For the models trained with $\ell_{2}$ perturbation, we evaluate with $\ell_{2}$ perturbation magnitude $\epsilon=128/255$. We also use Auto Attack with the same $\epsilon$ value as in PGD attack.
@@
Medical Application. Below, we detail the setup employed for our experiments on the medical datasets outlined in Section 5.4. For this experiment, we trained models using a ResNet-18 (He et al., 2016) architecture We consider $\ell_{\infty}$ perturbations with $\epsilon=0.1$. We set the attack step size to $0.02$, and the number of PGD steps to $10$. The final model is trained with SSAT using the Adam optimizer (Kingma, 2014) with a learning rate of $0.001$. We evaluate model robustness with PGD attacks with a step size of $0.02$ and $20$ attack steps.
Table 3. Comparisons of SSAT methods with different selection schemes based on average efficiency score and robust accuracy. Here, we compute the average of the results for $\alpha=10\%$ and $\alpha=20\%$ for our selection methods, and for the “No Selection” method, we use $\alpha=100\%$, taking the average of the external and generated data results for each setting.
[…]
SVHN CIFAR-10
Random $1.00$ $83.1\pm 1.04$ $56.8\pm 0.59$
PCS $0.89$ $84.0\pm 0.92$ $57.4\pm 0.53$
LCS-GMM $0.87$ $84.2\pm 0.89$ $57.8\pm 0.67$
LCS-KM $0.88$ $86.6\pm 0.33$ $59.3\pm 1.15$
No Selection $0.33$ $86.2\pm 0.15$ $62.2\pm 0.35$
Appendix C Detailed Comparisons of Different Data Selection Schemes
This section presents a comprehensive comparison of all the selection methods, including the case where no selection is applied. It highlights the tradeoff between performance and efficiency, offering a detailed analysis of the conditions under which each method is most suitable. Table 3 summarizes the performance of SSAT with different data selection strategies in terms of efficiency and robust accuracy. Each method strikes a balance between computational cost and accuracy improvements. Random selection serves as a baseline with minimal computation but lacks precision, while our strategic selection methods prioritize boundary points, which achieve progressively improved performance but with moderate increases in computational demands. Using $100\%$ of the data without selection yields the highest robust accuracy but at a significant computational cost. More specifically, we calculate a normalized efficiency score for each method defined as $\text{Eff Score}_{ij}=\frac{T_{\text{min},j}}{T_{ij}}$, where $T_{\text{min},j}$ is the minimum training time for dataset $j$, and $T_{ij}$ is the actual training time for method $i$. The overall efficiency score for method $i$ is the average across all datasets computed as $\frac{1}{n}\sum_{j}\text{Eff Score}_{ij}$, where $n$ denotes the number of datasets and sources considered. Effectiveness is measured by the average robust accuracy across CIFAR-10 and SVHN. In particular, we compute the average of the results for $\alpha=10\%$ and $\alpha=20\%$ for our selection methods, and for the no-selection method, we use $\alpha=100\%$. In all cases, we take the average of the results from both the external and generated data. These metrics are introduced to provide a fair comparison between methods.
Appendix D Additional Experiments
D.1. Other Hyperparameter Analysis Details and Results
In Section 6.2, we present a detailed analysis of how accuracies are influenced by different hyperparameter values. In these experiments, we varied one hyperparameter at a time while keeping the others constant. Figure LABEL:fig:betacifar illustrates how robust accuracy on the CIFAR-10 dataset changes with $\beta$ values when $20\%$ additional data is incorporated from the $1$M data samples generated by the DDPM model, using a per-batch ratio of $0.3$. Figure LABEL:fig:betasvhn shows the variation in robust accuracy for the SVHN dataset as $\beta$ changes, with $10\%$ drawn from the SVHN extra dataset and a per-batch ratio of $0.5$. In Figure LABEL:fig:ratiocifar, we present how robust accuracy for the CIFAR-10 dataset is affected by varying the ratio of per-batch ratio $\gamma$ when $20\%$ extra data from the $1$M generated DDPM samples is used, with $\beta$ fixed at $0.4$. Similarly, Figure LABEL:fig:ratiosvhn displays the effect of varying the per-batch ratio on robust accuracy for the SVHN dataset, using $10\%$ data from the SVHN extra dataset and fixing $\beta$ at $0.6$. In addition, Figures LABEL:fig:cofarper and LABEL:fig:svhnper explore how robust accuracy varies with the amount of data used. For CIFAR-10 in Figure LABEL:fig:cofarper, the per-batch ratio is set to $0.5$, and $\beta$ is fixed at $0.4$. For the SVHN dataset in Figure LABEL:fig:svhnper, the per-batch ratio is set to $0.5$, and $\beta$ is set to $0.6$. We aim to capture the robust accuracy at the optimal epoch in each case, and we observe that for Figures LABEL:fig:cofarper and LABEL:fig:svhnper, as the amount of data increases, the optimal accuracy is reached at a considerably later epoch.

### v2

@@ V-C Hyperparameter Sensitivity
Data Reduction Ratio $\bm{\alpha}$. Figures 4c and 4d describe how accuracies vary with the ratio parameter $\alpha$. We analyze how the quantity of extra unlabeled data influences the effectiveness of our data reduction method, which is designed to prioritize points near the decision boundary. When the available data is limited, our algorithm excels by selecting a higher proportion of boundary-adjacent data, achieving a balanced selection that enhances robust accuracy. However, as the volume of unlabeled data increases, the algorithm’s focus on boundary points can inadvertently lead to overfitting, diminishing the improvement gained on the model’s generalization ability. Interestingly, as the data volume continues to grow and more points are added, the performance begins to improve once again, suggesting that the initial overfitting is mitigated by the sheer abundance of data, leading to a recovery in accuracy.
(a) Fine-Tuning Epochs $S$
(b) Regularization Strength $\lambda$
Figure 5: Clean and robust accuracy curves of SSAT with LCG-KM using generated data on CIFAR-10 under $\alpha=10\%$ by varying: (a) the number of fine-tuning epochs $S$, and (b) the regularization strength parameter $\lambda$.
Guided DDPM Finetuning. We conduct ablation studies on our guided DDPM fine-tuning approach to better illustrate its behavior. In particular, we focus on two key hyperparameters, the number of fine-tuning epochs $S$ and the regularization strength parameter $\lambda$, and study their impact on accuracy and robustness performance when training SSAT algorithms on CIFAR-10 with LCG-KM and $\alpha=10\%$. We also investigate the quality of the data samples generated by our fine-tuned DDPM using both quantitative and qualitative metrics.
Figure 5a illustrates our ablation results on $S$. We vary the number of fine-tuning epochs $S\in\{5,15,25,35,45\}$ with $\lambda=0.5$, and report the final clean and robust accuracies after the whole SSAT process. Figure 6 visualizes the images generated by our DDPM models with different fine-tuning epochs. We observe that choosing a small number of epochs for fine-tuning DDPM (e.g., 10102020 epochs) is usually sufficient to achieve the best robustness performance, likely because the starting DDPM is already pre-trained for CIFAR-10. Training for too many epochs can lead to a degradation in image quality and loss of diversity, thereby adversely influencing the final robust accuracies. In addition, Figure 5b shows the results for the regularization strength parameter $\lambda\in\{0.1,0.3,0.5,0.7\}$ with fixed epoch number $S=15$. Our findings show that choosing $\lambda$ within the range of $[0.4,0.6]$ results in the best performance, striking a balance between the quality of the generated images and their proximity to the model’s decision boundary.
Figure 6: Visualization of generated data across different fine-tuning epochs
VI Conclusion and Future Work
In this work, we proposed various unlabeled data reduction schemes to improve the efficiency of SSAT while preserving robust accuracy. In particular, we emphasized that strategically selecting or generating more boundary-adjacent data points can significantly improve the efficiency of semi-supervised adversarial training algorithms, without losing their robustness advantages. Among our proposed methods, we found that latent-based approaches using k-means clustering, including LCS-KM and LCG-KM, perform the best, as they reduce computational costs and memory requirements.
Limitation & Future Work. We acknowledge that our work has several limitations. While the selection scheme with LCS-KM consistently achieves the best performance, its success requires carefully choosing a few hyperparameters (e.g., to avoid overfitting). Future studies can leverage AutoML algorithms to automatically determine these hyperparameters. Additionally, it remains unclear whether our best-performing latent-clustering-based methods (i.e., LCS-KM and LCG-KM) can be further optimized to achieve better performance. Theoretical analysis of how to solve the optimization problems (Equations 4 and 6) in a more principled way would be helpful. Furthermore, our guided DDPM fine-tuning approach proposes to adapt a pre-trained model for boundary-seeking data generation. An interesting future direction is to investigate whether training-free sampling-based guidance [34] or more advanced generative models [35, 36] can enhance the efficiency and robustness of SSAT. Despite these limitations, we believe our work can serve as an important initial step towards the more efficient and scalable development of robust models.
Availability
The implementation of the latent-based data reduction methods and all our experiments are available as open-source code at: https://github.com/TrustMLRG/EfficientSSAT.
Appendix A Related Work
A-A Adversarially Robsut Learning
Adversarial examples are inputs crafted with small perturbations that are designed to mislead model predictions [1]. The prevalence of adversarial examples with deep neural networks poses a critical challenge in modern machine learning, especially for security-critical or safety-critical applications. Earlier-proposed attacks like fast gradient sign method (FGSM) [2] and projected gradient descent (PGD) [7] introduced gradient-based techniques to find adversarial examples using constrained optimization, while Carlini and Wagner proposed unconstrained CW attacks [37], which search for the minimal perturbation that causes misclassification. Recent attacks like Auto-PGD [38] automate gradient steps for higher efficiency in finding worst-case perturbations, while AutoAttack [17] is an ensemble of diverse parameter-free attacks, making them more effective in evaluating model robustness.
Adversarial Training. On the defense side, adversarial training and its variants are the most popular methods for achieving robustness. For instance, Madry et al. proposed to train DNNs on adversarially perturbed inputs using PGD attacks [7], while Zhang et al. introduced TRADES [9], which balances the trade-off between robustness and standard accuracy. Nevertheless, adversarial training algorithms are often criticized for overfitting adversarially perturbed samples produced during training time and for their high computation costs, which hinder their deployment in real-world applications. In particular, Rice et al. illustrated through comprehensive experiments that the robust overfitting phenomenon is prevalent in adversarial training [28], where early stopping is shown to be effective in alleviating such an issue. Techniques, such as Free AT [18] and Fast AT [19], offer promising solutions to reduce the computation for adversarial training. Nevertheless, they suffer from decreased robustness performance.
Moreover, a line of research pointed out that treating all the data points equally in adversarial training is not optimal for robust learning since different examples can impact model robustness differently. For example, Zhang et al. argued that data points far from decision boundaries are inherently more secure and less likely to be influential in robust learning against adversarial perturbations [16]. As a result, these points should be given less weight during training to prevent the model from overfitting them, which could negatively affect its generalization on more vulnerable or unseen data. In a similar vein, Hua et al. advocated for a more targeted approach [24], recommending that PGD training be applied primarily to examples situated near the decision boundaries. This strategy focuses the model’s attention on the points that are most likely to enhance its robustness, as adversarial perturbations on boundary-adjacent data are more likely to result in meaningful improvements to the model’s resistance to attacks. Our work builds on these ideas by strategically identifying and prioritizing boundary-adjacent points from the unlabeled dataset to address the unique challenges associated with self-supervised adversarial training methods.
Self-Supervised Adversarial Training. Utilizing unlabeled data to improve model performance has become an active and rapidly advancing field of research [39]. In the adversarial context, Schmidt et al. [10] highlighted a key observation: the sample complexity required for robust learning far exceeds that of standard learning. They showed that achieving adversarially robust generalization often necessitates a much larger dataset than traditional learning approaches, posing a challenge when labeled data is scarce or costly. Subsequently, a line of works proposed to boost model robustness by involving additional unlabeled data, often acquired from a similar but slightly different distribution, using self-supervised methods in adversarial training [11, 40, 41, 12]. In particular, Carmon et al. proposed robust self-training methods to address the sample complexity issue [11], which first trains an intermediate model using available labeled data and then leverages the model to generate pseudo-labels for unlabeled data. The pseudo-labeled samples, combined with the original labeled data, are subsequently used to train a final robust model. Concurrently, Najafi et al. proposed an optimization framework incorporating labeled and unlabeled data [41], while providing formal guarantees on adversarial robustness.
Leveraging synthetic data produced by state-of-the-art generative models to build robust models has also been explored extensively in recent literature [13, 14, 15]. For instance, Gowal et al. employed various unconditional generative models [13], including variational autoencoders (VAEs), generative adversarial networks (GANs), and the more advanced denoising diffusion probabilistic models (DDPMs), to produce a large synthetic dataset to be incorporated in adversarial training, which improves robustness further. While all the aforementioned SSAT methods enrich the training dataset, they require considerable additional unlabeled data to ensure effective robustness enhancement. In this work, we explore data selection schemes to reduce the amount of unlabeled or generated data that are both efficient and effective for self-supervised adversarial training to achieve good robustness performance.
A-B Data-Efficient Deep Learning
Dataset Distillation. Dataset distillation aims to distill large datasets into compact, synthetic, yet informative datasets by matching gradients or optimizing specific loss functions, allowing models to achieve comparable performance with significantly fewer data points. For instance, Wang et al. proposed a foundational approach to dataset distillation [42], proposing methods to condense large datasets into smaller, synthetic subsets while preserving model performance. Later, Zhao et al. introduced the concept of gradient matching [43], where distilled data points are optimized to match the gradients of the original dataset. This technique largely improves the distillation performance, enabling models to learn more effectively from reduced data. Additionally, Sucholutsky and Schonlau proposed soft-label dataset distillation, which uses probabilistic labels for distilled data [44], facilitating the training of complex models with minimal data while maintaining strong performance. Despite its success in standard deep learning, dataset distillation is often impractical for robust learning due to its high computational demands. Moreover, it assumes that all the examples in the entire dataset are equally important for distillation, ignoring the fact that different data points may have varying influences on model performance.
Coreset Selection. On the other hand, coreset selection [45, 46] aims to identify a small, representative data subset that retains the model performance when trained on this reduced dataset. These approaches often rely on diversity-based metrics or optimization techniques to ensure that the selected subset preserves sufficient information for effective learning. For example, Kaushal et al. utilized diverse models to select training data subsets to reduce labeling efforts [20], while Xia et al. introduced the concept of “moderate coreset” [21], focusing on data points with scores near the median to construct subsets that generalize well across different scenarios based on score distributions. Similarly, Mirzasoleiman et al. proposed selecting a weighted subset of training data that approximates the full gradient by maximizing a submodular function [22].
In the adversarial context, Dolatabadi et al. proposed adversarial coreset selection [47], offering a task-specific solution to enhance the computational efficiency of adversarial training. In particular, the method constructs a compact data subset by minimizing the gradient approximation error. The constructed subset is then utilized for adversarial training, with selection taking place at regular intervals. While effective in reducing computational overhead for vanilla adversarial training, the growing reliance on large volumes of generated or external data—such as in self-supervised adversarial training—introduces new challenges. These approaches often expand the dataset size to improve robust accuracy, eventually increasing the total training time. In such settings, repeatedly selecting data during training may become computationally infeasible. In summary, our work complements the above-mentioned methods for efficient deep learning, focusing on developing more data selection schemes without compromising robustness or scalability, particularly for SSAT when large-scale unlabeled data are involved.
@@ Appendix B Detailed Experimental Setup
SVHN. SVHN is naturally divided into a core training set comprising approximately 7373K images and an extra training set of around 531531K images. Initially, the model is trained on the 7373K labeled images. To evaluate the generalizability of our selection schemes, we apply it to a synthetic dataset generated using a DDPM model [13].
@@
Configuration. For model training, we adopt a WideResNet 16-8 (WRN-16-8) architecture [48] in all our SVHN experiments. We generate adversarial examples using PGD attacks exactly as implemented by Zhang et al. [9], with step size $0.007$, $10$-step PGD attack iterations, and $\ell_{\infty}$ perturbation magnitude $\epsilon=0.015$. Hyperparameters are set the same as in Carmon et al. [11] except for the number of epochs. For SVHN, we use a training batch size of $128$. All of our experiments that utilize all of the extra 531K data are run for $200$ epochs and the experiments with 11M generated data are run for $400$ epochs. To prevent overfitting and reduce computational complexity, we run our experiments with selected data for $75$ epochs with early stopping. Following prior literature [11], we adopt an SGD optimizer with a weight decay factor of $5\cdot 10^{4}$ and an initial learning rate of $0.1$ with a cosine learning rate annealing to train the model.
@@
Evaluation. For the attack evaluation to calculate the robust accuracy using PGD attack, we keep the parameters similar to that of Carmon et al. [11] for better comparison. We use step size $0.005$, number of attack steps $K=100$. We evaluate models at $\epsilon=0.015$, which is the same as the value we used during training. We also use Auto Attack with the same $\epsilon$ value as in PGD attack.
@@
CIFAR-10. The CIFAR-10 dataset has 50K labeled images. We use the $80$ Million Tiny Images (80M-TI) dataset, of which CIFAR-10 is a manually labeled subset, to obtain extra data. However, most images in 8080M-TI do not correspond to CIFAR-10 image categories. Carmon et al. [11] used an 11-way classifier to distinguish CIFAR-10 classes and an 11th “non-CIFAR-10” class using a WideResNet 28-10 model. Each class selected an additional 5050K images from 80M-TI using the model’s predicted scores to create a 500500K pseudo-labeled dataset, which we used in our experiments. We train the intermediate model with 5050K labeled data and use this model to select data from the 500500K pseudo-labeled data or a synthetic dataset. We perform the same experiments as in SVHN, except that we do not have the ground-truth labels of the additional unlabeled dataset. We also conducted further experiments utilizing 11M synthetical data generated by the DDPM model [13]. For our fine-tuning experiments, we start with a publicly available pretrained DDPM model [49] and further fine-tune it for a few epochs to generate the desired images, which are subsequently used for our final training.
@@
Configuration. For our main CIFAR-10 experiments, we use a WideResNet 28-10 (WRN-28-10) [48] architecture to train models. Similar to SVHN experiments, we set step size $0.007$, PGD attack iterations as $10$ and $\ell_{\infty}$ perturbation magnitude $\epsilon=0.031$. For our adversarial training with $\ell_{2}$ perturbation, we use $\epsilon=128/255$. In section D-A when $\epsilon=4/255$, we use $0.0035$ as the PGD step size. For $\epsilon=16/255$, we use $0.015$ as the PGD step size. Hyperparameters used are the same as in Carmon et al. [11] except for the number of epochs. Here, we use a training batch size of $256$. The experiments using all the 500500K pseudo-labeled data are run for $200$ epochs, and using 1M generated data are run for $400$ epochs. For our experiments with limited data, we get the best results at $100$ epochs where we early stop the training process. Following prior literature [11], we adopt an SGD optimizer with a weight decay factor of $5\cdot 10^{4}$ and an initial learning rate of $0.1$ with a cosine learning rate annealing to train the model.
@@
Evaluation. The PGD attack evaluation is conducted similarly to that of Carmon et al. [11] for fair comparisons. We use step size $\alpha=0.01$, number of attack steps $K=40$. We consider $\ell_{\infty}$ perturbations with $\epsilon=0.031$, the same as the perturbation magnitude we used during training. For the models trained with $\ell_{2}$ perturbation, we evaluate with $\ell_{2}$ perturbation magnitude $\epsilon=128/255$. We also use Auto Attack with the same $\epsilon$ value as in PGD attack.
@@
Medical Application. Below, we detail the setup employed for our experiments on the medical datasets outlined in Section V-A. For this experiment, we trained models using a ResNet-18 [50] architecture We consider $\ell_{\infty}$ perturbations with $\epsilon=0.1$. We set the attack step size to $0.02$, and the number of PGD steps to $10$. The final model is trained with SSAT using the Adam optimizer [51] with a learning rate of $0.001$. We evaluate model robustness with PGD attacks with a step size of $0.02$ and $20$ attack steps.
Appendix C Additional Experiments
(a) CIFAR-10
(b) SVHN
Figure 7: Illustration of clean and robust accuracy curves of SSAT algorithms with LCS-KM by varying per-batch ratio $\gamma$, which defines how much labeled and unlabeled data are included in each mini-batch, on: (a) CIFAR-10 with $20\%$ extra unlabeled data selected from 11M generated images using DDPM, and (b) SVHN with $10\%$ selected from the external 531531K SVHN images.